OpenStudy (anonymous):

Verify tan(x + pi/2) = -cotx

OpenStudy (anonymous):

tan ( x+ pi/2) = tan ( 90 + x) since the angle is in the second quadrant .where tan is negative. thus the negative sign. and the angle from the horizontal is 180-(90+x) =90-x..hence it changes to cot therefore tan(90=x) = - cotx

OpenStudy (anonymous):

thanks

OpenStudy (shubhamsrg):

To cut out on the theory, mathematically you can prove it like this : tan(pi/2 +x) = tan(pi/2 - (-x)) = cot(-x) = -cotx

OpenStudy (anonymous):

i think u need some other method, the very basic method ?